OpenStudy (anonymous):

some please work this problem completely step by step show ur work so i can follow along divide & simplify z^2-49 divided by 2z-14 ------- --------- z^2-8z+16 z^2-2z-63

OpenStudy (anonymous):

To divide fractions, you always have to invert (flip) the second fraction and multiply

OpenStudy (anonymous):

\[\frac{z^2-49}{z^2-8z+16} * \frac{z^2-2z-63}{2z-14}\]

OpenStudy (anonymous):

Next factor each part.

OpenStudy (anonymous):

\[\frac{(z-7)(z+7)}{(z-4)(z-4)} * \frac {(z-9)(z+7)}{2(z-7)}\]

OpenStudy (anonymous):

sorry I'm so slow typing those equations

OpenStudy (anonymous):

Then see what cancels (anything in the top with anything in the denominator)

OpenStudy (anonymous):

so the z-7 will cancel. Nothing else will cancel

OpenStudy (anonymous):

Everything that is still there is your answer....\[\frac{(z-9)(z+7)^2}{2(z-4)^2}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes